### v2

# There Was Never a Bottleneck in Concept Bottleneck Models

Abstract
Deep learning representations are often difficult to interpret, which can hinder their deployment in sensitive applications. Concept Bottleneck Models (CBMs) have emerged as a promising approach to mitigate this issue by learning representations that support target task performance while ensuring that each component predicts a concrete concept from a predefined set. In this work, we argue that CBMs do not impose a true bottleneck: the fact that a component can predict a concept does not guarantee that it encodes only information about that concept. This shortcoming raises concerns regarding interpretability and the validity of intervention procedures. To overcome this limitation, we propose Minimal Concept Bottleneck Models (MCBMs), which incorporate an Information Bottleneck (IB) objective to constrain each representation component to retain only the information relevant to its corresponding concept. This IB is implemented via a variational regularization term added to the training loss. As a result, MCBMs yield more interpretable representations, support principled concept-level interventions, and remain consistent with probability-theoretic foundations.
1 Introduction
Most machine learning models operate by learning data representations—compressed versions of the input that retain the essential information needed to solve a given task [1]. However, these representations often encode information in ways that are not easily interpretable by humans. This lack of interpretability becomes especially problematic in sensitive domains such as healthcare [2, 3, 4], finance [5, 6], and autonomous driving [7, 8]. To address this issue, Concept Bottleneck Models (CBMs) have been proposed, which enforce representations to be defined in terms of a set of human-understandable concepts [9].
Formally, given a task $\bm{y}$ and a set of concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}$, Vanilla Models (VMs) are trained with a single objective: their representations $\bm{z}$ should capture the information necessary to predict $\bm{y}$ accurately. CBMs extend this framework by adding a second objective: each concept $c_{j}$ must be recoverable from a designated component $z_{j}\in\bm{z}$. By enforcing this additional constraint, CBMs are claimed to offer: (i) improved interpretability of the representation space, and (ii) the ability to intervene on specific concepts by modifying $z_{j}$ and propagating these changes to the predictions.
However, CBMs are prone to a phenomenon known as information leakage [10, 11], where the representation $\bm{z}$ encodes input information that cannot be attributed to the predefined concepts $\bm{c}$. We refer to this additional information as nuisances $\bm{n}$. Information leakage raises two main concerns: (i) it undermines interpretability, since $z_{j}$ cannot be fully explained by its corresponding concept $c_{j}$; and (ii) it compromises the validity of interventions—modifying $z_{j}$ may alter not only the associated concept $c_{j}$, but also other unintended information encoded in $z_{j}$.
We argue that information leakage stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) [12] that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$. While the second objective in CBMs encourages each $z_{j}$ to retain $c_{j}$ in its entirety, it does not enforce that $z_{j}$ captures only information about $c_{j}$. In the worst-case scenario, $z_{j}$ could encode the entire input $\bm{x}$ and still satisfy this objective.
Table 1: Differences between VMs, CBMs and MCBMs.
VMs CBMs MCBMs
Does $z_{j}$ encode all $c_{j}$?
Does $z_{j}$ encode only $c_{j}$?
To address this issue, we propose Minimal Concept Bottleneck Models (MCBMs), which incorporate an IB into each $z_{j}$. This ensures that $z_{j}$ not only retains all the information about its associated concept $c_{j}$, but also excludes any information unrelated to $c_{j}$, as summarized in Table 1. The name reflects that $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ [13, 14], in contrast to traditional CBMs where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$. As illustrated in Figure 1, this design yields disentangled representations that directly address the two shortcomings of CBMs: (i) it improves interpretability, since $z_{j}$ can be fully explained by its corresponding concept $c_{j}$; and (ii) it enables valid interventions—modifying $z_{j}$ affects only the associated concept $c_{j}$.
In Section 2, we connect the data generative process to MCBMs through information-theoretic quantities, showing that the IB can be implemented via a variational loss. In Section 3, we review alternative approaches to address information leakage. In Section 4, we present experiments demonstrating that MCBMs enforce a true bottleneck, thereby enhancing interpretability and intervenability compared to existing alternatives. Finally, in Section 5, we show that the assumptions made in CBMs to enable interventions are theoretically flawed.
Figure 1: In Vanilla Models, concepts and nuisances may be arbitrarily entangled in the representation, and a variable $z_{j}$ may capture only part of a concept (depicted as paler colors). In CBMs, each $z_{j}$ encodes all information about its corresponding concept $c_{j}$, but may also capture some information about nuisances (e.g., $z_{1}$) or other concepts (e.g., $z_{2}$). In contrast, MCBMs enforce that each representation variable $z_{j}$ encodes all—and only—the information about its corresponding concept.
2 From Data Generative Process to MCBMs
2.1 Data Generative Process
For this scenario, we consider inputs $\bm{x}\in\mathcal{X}$, targets $y\in\mathcal{Y}$, concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}\in\mathcal{C}$ and nuisances $\bm{n}\in\mathcal{N}$, such that $p(\bm{x},\bm{y},\bm{c},\bm{n})=p(\bm{x}|\bm{c},\bm{n})p(\bm{y}|\bm{x})$, i.e., the inputs $\bm{x}$ are described by the concepts $\bm{c}$ and the nuisances $\bm{n}$, and the targets $\bm{y}$ are fully described by the input $\bm{x}$. The only difference between $\bm{c}$ and $\bm{n}$ is that the former are observed, while the latter are not or, in other words, labels on $\bm{c}$ are provided to us. We assume access to a training set $\{\bm{x}^{(i)},\bm{y}^{(i)},\bm{c}^{(i)}\}_{i=1}^{N}$, which defines the empirical distribution $p(\bm{x},\bm{y},\bm{c})=\sum_{i=1}^{N}\delta\left(\bm{x}-\bm{x}^{(i)}\right)\delta\left(\bm{y}-\bm{y}^{(i)}\right)\delta\left(\bm{c}-\bm{c}^{(i)}\right)$. The graphical model corresponding to this generative process is shown in Figure 2(a) for the case of two concepts.
2.2 Vanilla Models
In machine learning, the most commonly studied problem is that of predicting $\bm{y}$ from $\bm{x}$, which serves as a foundation for more specialized tasks. We refer to models trained to address this problem as Vanilla Models (VMs). These models typically operate by first extracting an intermediate representation $\bm{z}\in\mathcal{Z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$. Subsequently, a prediction $\hat{y}\in\mathcal{Y}$ is produced from $\bm{z}$ using a task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$. Since $\bm{z}$ is intended to facilitate accurate prediction of $\bm{y}$, the mutual information between $\bm{z}$ and $\bm{y}$, denoted $I(Z;Y)$, should be maximized. In Appendix B.1, we formally show that:
$\max_{Z}I(Z;Y)=\max_{\theta,\phi}\mathbb{E}_{p(\bm{x},\bm{y})}\left[\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\hat{\bm{y}}|\bm{z})\right]\right]$ (1)
Figure 2(b) shows the graphical model of a Vanilla Model with a two-dimensional representation $\bm{z}$. Black edges represent the encoder $p_{\theta}(\bm{z}\mid\bm{x})$, while green edges indicate the task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$. The encoder is typically chosen to be deterministic, i.e., $p_{\theta}(\bm{z}\mid\bm{x})=\delta\left(\bm{z}-f_{\theta}(\bm{x})\right)$. However, for tractability reasons (see Section 2.4), we adopt a stochastic formulation where $p_{\theta}(\bm{z}\mid\bm{x})=\mathcal{N}\left(\bm{z};f_{\theta}(\bm{x}),\sigma_{x}^{2}I\right)$, as summarized in Table 2. The choice of task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ depends on the structure of the output space $\mathcal{Y}$, also detailed in Table 2. The objective in Equation 1 corresponds to minimizing the cross-entropy loss when $\bm{y}$ is binary or multiclass, and to minimizing the mean squared error between $\bm{y}$ and $g^{y}_{\phi}(z)$ when $\bm{y}$ is continuous.
2.3 Concept Bottleneck Models
Vanilla Models generally lack interpretability with respect to the known concepts $\bm{c}$, as the encoder $f_{\theta}$ is often opaque and difficult to analyze. Moreover, these models tend to entangle the concepts in such a way that it becomes intractable to determine how individual concepts influence specific components of the latent representation $\bm{z}$, and consequently the predictions $\hat{y}$. Concept Bottleneck Models (CBMs) have been introduced to address this limitation. In a CBM, each concept $c_{j}$ is predicted from a dedicated latent representation $z_{j}$ via a concept head $q(\hat{c}_{j}\mid z_{j})$. Consequently, $z_{j}$ must encode all information about $c_{j}$—that is, $z_{j}$ must be a sufficient representation for $c_{j}$. This requirement can be formalized as maximizing the mutual information $I(Z_{j};C_{j})$, for which the following identity—proved in Appendix B.2—is employed:
$\max_{Z_{j}}I(Z_{j};C_{j})=\max_{\theta,\phi}\mathbb{E}_{p(\bm{x},c_{j})}\left[\mathbb{E}_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\right]$ (2)
As illustrated in Figure 2(c), CBMs extend Vanilla Models by incorporating a concept head $q(\hat{c}_{j}\mid z_{j})$, depicted with blue arrows. These models jointly optimize the objectives in Equations 1 and 2. As detailed in Table 2, the form of the concept head $q_{\phi}(\hat{c}_{j}\mid z_{j})$ depends on the nature of the concept space $\mathcal{C}$. The objective in Equation 2 corresponds to minimizing the cross-entropy loss when $c_{j}$ is binary or multiclass, and the mean squared error between $c_{j}$ and $g_{\phi}^{c}(z_{j})$ when $c_{j}$ is continuous.
How are Interventions Performed in CBMs?
As explained in Section 1, a key advantage often attributed to CBMs is their ability to support concept-level interventions. Suppose we aim to estimate $p(\hat{\bm{y}}\mid c_{j}=\alpha,\bm{x})$. In CBMs, this intervention is performed through the latent representation $z_{j}$:
[…]
However, the conditional distribution $p(z_{j}\mid c_{j})$ is not defined—there is no directed path from $c_{j}$ to $z_{j}$ in Figure 2(c). Intuitively, because $z_{j}$ may encode information about $\bm{x}$ beyond $c_{j}$, it cannot be fully determined by $c_{j}$ alone. This raises a key question: how can interventions be performed in CBMs if $p(z_{j}\mid c_{j})$ is unknown? To make interventions feasible, CBMs typically impose two constraints:
  1. (i)
    Concepts $c_{j}\in\bm{c}$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest [15]).
  2. (ii)
    The concept head is defined as $q_{\phi}(c_{j}\mid z_{j})=\sigma(z_{j})$, where $\sigma$ denotes the sigmoid function.
Since $\sigma$ is invertible, this setup permits defining $p(z_{j}\mid c_{j})\approx\sigma^{-1}(c_{j})$. However, this is ill-defined at the binary extremes, as $\sigma^{-1}(1)=-\sigma^{-1}(0)=\infty$. To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are set as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively [9]. This workaround, however, introduces two crucial issues discussed in Section 5.
Table 2: Distributions considered in this work. $f_{\theta}$ is typically modeled by a large neural network while $g^{y}_{\phi}$, $g^{c}_{\phi}$ and $g^{z}_{\phi}$ by simpler neural networks. We consider $\bm{z}$ to be always continuous.
$p_{\theta}(z|x)$ $q_{\phi}(\hat{y}|z)$ $q_{\phi}(\hat{c}_{j}|z_{j})$ $q_{\phi}(\hat{z}_{j}|c_{j})$
Binary - $\text{Bernoulli}\left(g^{y}_{\phi}(z)\right)$ $\text{Bernoulli}\left(g^{c}_{\phi}(z_{j})\right)$ -
Multiclass - $\text{Categoric}\left(g^{y}_{\phi}(z)\right)$ $\text{Categoric}\left(g^{c}_{\phi}(z_{j})\right)$ -
Continuous $\mathcal{N}\left(f_{\theta}(x),\sigma^{2}_{x}I\right)$ $\mathcal{N}\left(g^{y}_{\phi}(z),\sigma_{\hat{y}}^{2}I\right)$ $\mathcal{N}\left(g^{c}_{\phi}(z_{j}),\sigma_{\hat{c}}^{2}I\right)$ $\mathcal{N}\left(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I\right)$
2.4 Minimal Concept Bottleneck Models
As discussed in Section 1, CBMs lack an explicit mechanism for enforcing a bottleneck, which often undermines their intended advantages. To address this limitation, we introduce Minimal Concept Bottleneck Models (MCBMs), which explicitly impose an Information Bottleneck. This ensures that each $z_{j}$ retains all information about its associated concept $c_{j}$, while excluding information unrelated to $c_{j}$. In other words, $z_{j}$ becomes a minimal sufficient representation of $c_{j}$. To achieve this, we introduce a representation head $q_{\phi}(\hat{z}_{j}\mid c_{j})$ that predicts $z_{j}$ from $c_{j}$. This encourages $z_{j}$ to discard information unrelated to $c_{j}$, as doing so improves the predictive accuracy of $q_{\phi}(\hat{z}_{j}\mid c_{j})$. Formally, this objective corresponds to minimizing the conditional mutual information $I(Z_{j};X\mid C_{j})$. We note that if $z_{j}$ is a representation of $\bm{x}$, then the following propositions are equivalent: (i) $I(Z_{j};X|C_{j})=0$, (ii) the Markov Chain $X\leftrightarrow C_{j}\leftrightarrow Z_{j}$ is satisfied and (iii) $p(z_{j}|c_{j})=p(z_{j}|x)$. To minimize $I(Z_{j};X\mid C_{j})$, we leverage the following identity, which is proven in Appendix B.3:
$\min_{Z_{j}}I(Z_{j};X|C_{j})=\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta}(z_{j}|x)||q_{\phi}(\hat{z}_{j}|c_{j})\right)\right]$ (4)
Figure 2(d) illustrates how MCBMs extend CBMs by introducing the representation head $q(\hat{z}_{j}\mid c_{j})$, depicted with red arrows. These models are trained to jointly optimize the objectives given in Equations 1, 2, and 4. Although the KL Divergence in Equation 4 does not admit a closed-form solution in general, we show in Appendix B.4 that, under the distributional assumptions listed in Table 2, it reduces to the mean squared error between $f_{\theta}(x)$ and $g^{z}_{\phi}(c_{j})$.
How are Interventions Performed in MCBMs?
In contrast to CBMs, MCBMs explicitly constrain $z_{j}$ to contain only information about $c_{j}$. As a result, modifying $z_{j}$ corresponds to intervening solely on $c_{j}$. This is reflected in Figure 2(d), where MCBMs introduce a directed path from $c_{j}$ to $z_{j}$ through the intermediate variable $\hat{z}_{j}$. This structure permits the computation of:
$p(z_{j}|c_{j})=\int p(z_{j}|\hat{z}_{j})q_{\phi}(\hat{z}_{j}|c_{j})\,d\hat{z}_{j}$ (5)
which simplifies to $p(z_{j}\mid c_{j})=q_{\phi}(z_{j}\mid c_{j})$ when $p(z_{j}\mid\hat{z}_{j})=\delta(z_{j}-\hat{z}_{j})$, i.e., once the objective in Equation 4 is optimized and $z_{j}$ encodes exclusively $c_{j}$. As shown in Table 2, we define the encoder distribution as $q_{\phi}(z_{j}\mid c_{j})=\mathcal{N}(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I)$, where mean function $g^{z}_{\phi}(c_{j})$ is chosen according to the following rules:
  1. (i)
    For binary concepts, $g^{z}_{\phi}(c_{j})=\lambda$ if $c_{j}=1$, and $g^{z}_{\phi}(c_{j})=-\lambda$ otherwise.
  2. (ii)
    For categorical concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot\mathrm{one\_hot}(c_{j})$. This mirrors the Prototypical Learning [16] approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed.
  3. (iii)
    For continuous concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot c_{j}$.
Here, $\lambda$ is a scaling constant that controls the norm of the latent representation, fixed to $\lambda=3$ in all experiments. Regarding the variance term, we set: (i) $\sigma_{x}=0$in the case of CBMs to obtain a deterministic encoder in line with their original formulation, and (ii) $\sigma_{x}=\sigma_{\hat{z}}=1$for MCBMs.
Practical Considerations for Optimizing MCBMs
To optimize MCBMs, we combine the three previously introduced objectives, incorporating two key considerations. First, we approximate the expectations over $p(\bm{x},\bm{y})$ and $p(\bm{x},c_{j})$ using the empirical data distribution, replacing integrals with summations over the dataset. Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick [17]: $\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big|f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 2), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$. Combining the considerations above yields the final objective for training MCBMs, as shown in Equation 6, where $\beta$ and $\gamma$ are hyperparameters. The first term corresponds to the objective used in Vanilla Models, the second term is introduced in CBMs, and the third is specific to MCBMs. Detailed training algorithms for the various cases listed in Table 2 are provided in Appendix C.
$\begin{split}\max_{\theta,\phi}\sum_{k=1}^{N}\sum_{i}\log q_{\phi}\left(\hat{\bm{y}}\big|f^{\prime}_{\theta}\left(x^{(k)},\epsilon^{(i)}\right)\right)+&\beta\sum_{j=1}^{n}\log q_{\phi}\left(\hat{c}_{j}\big|f^{\prime}_{\theta,j}\left(\bm{x}^{(k)},\epsilon^{(i)}\right)\right)\\ &-\gamma\sum_{j=1}^{n}D_{KL}\left(p_{\theta}\left(z_{j}|\bm{x}^{(k)}\right)\big|\big|\ q_{\phi}\left(\hat{z}_{j}|c_{j}^{(k)}\right)\right)\end{split}$ (6)
(a) Data Generative Process
(b) Vanilla Model (VM)
(c) Concept Bottleneck Model (CBM)
(d) Minimal Concept Bottleneck Model (MCBM)
Figure 2: Graphical models of the different systems described for two concepts and two-dimensional representations. We provide in Appendix A the analogous figure for $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $\color[rgb]{0,1,1}p_{\theta}(\bm{z}|\bm{x})$ (cyan arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}|\bm{z})$ (green arrows). Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q(z_{j}|c_{j})$ (red arrows).
3 Related Work
Information Leakage in Concept Bottleneck Models
Information leakage occurs when the learned representation $\bm{z}$ encodes information outside the concept set $\bm{c}$, reducing both interpretability and intervenability [10, 11]. This arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios [18]. Concept Embedding Models (CEMs) [19] were proposed to mitigate the accuracy–interpretability trade-off in CBMs. However, this trade-off is fundamentally limited by the chosen concept set, and CEMs may even be less interpretable than standard CBMs, as their more entropic representations tend to amplify information leakage. Recently, this critique has been further formalized: Parisini et al. [20] proposed information-theoretic benchmarks for assessing interpretability and Makonnen et al. [21] introduced new metrics grounded in usable mutual information [22]. Havasi et al. [18] also introduced Hard Concept Bottleneck Models (HCBMs), which predict $\bm{y}$ from binarized concept predictions $\hat{c}_{j}$ rather than from $\bm{z}$ (see Appendix D), thereby imposing an ad-hoc Information Bottleneck. Extensions such as Autoregressive CBMs and Stochastic CBMs [18, 23] incorporate dependencies between concepts.
Information Bottleneck in Representation Learning
The Information Bottleneck (IB) [12] provides a principled way to balance preserving information about a factor with compressing the representation. In this framework, a representation is sufficient if it retains all the information about the variable of interest, and minimal if it contains only that information [24, 25, 26]. Computation of these quantities is intractable, so variational methods have been developed to derive tractable evidence bounds [27, 28].
4 Experiments
In this section, we present a series of experiments designed to empirically demonstrate that CBMs fail to impose an effective bottleneck, even in simple settings. We also examine the consequences of this limitation. In contrast, we show that MCBMs successfully enforce a bottleneck, thereby mitigating the issues that arise in its absence. Further implementation details—including encoder architectures and training hyperparameters for each experiment—are provided in Appendix E and at https://github.com/antonioalmudevar/minimal_cbm.
4.1 Do CBMs and MCBMs Leak Information?
(a) Data
(b) Representation
Figure 3: Some nuisances $\bm{n}_{y}\in\bm{n}$ affect the task $\bm{y}$ while others $\bm{n}_{\bar{y}}\in n$ do not. None of them should affect the representation $\bm{z}$ since it must be fully described by the concepts $\bm{c}$.
Some works assume that the task $\bm{y}$ is fully defined by the concepts $\bm{c}$. However, it is unrealistic to expect a finite set of human-understandable concepts to completely describe an arbitrarily complex task. In practice, certain nuisances $\bm{n}$ that describe the input $\bm{x}$ also influence $\bm{y}$. Specifically, we decompose the nuisances as $\bm{n}=\{\bm{n}_{y},\bm{n}_{\bar{y}}\}$, where $\bm{n}_{y}$ captures nuisances that, together with the concepts, describe $\bm{y}$, while $\bm{n}_{\bar{y}}$ comprises those independent of $\bm{y}$ (see Figure 3(a)). As discussed throughout this work, CBMs provide no incentive to remove information unrelated to the concepts. Moreover, since the objective of most models is to solve $\bm{y}$, they are incentivized to preserve not only $\bm{c}$ but also $\bm{n}_{y}$. While this may improve task performance, it comes at the expense of interpretability and valid interventions. For instance, if $z_{j}\in\bm{z}$ is intended to represent the concept $c_{j}\in\bm{c}$, , one might assume that modifying $z_{j}$ corresponds solely to intervening on $c_{j}$. However, if $z_{j}$ also encodes information about $\bm{n}_{y}$ (or even $\bm{n}_{\bar{y}}$ as we show later), then modifying $z_{j}$ also affects the nuisances, invalidating any causal conclusions. By introducing the Information Bottleneck in Equation 4, we explicitly constrain the model to remove nuisances from $z_{j}$, thereby restoring the validity of causal analyses involving $c_{j}$. We note, however, that this necessarily reduces task performance: if $\bm{c}$ is incomplete and solving $\bm{y}$ requires information from $\bm{n}$ (i.e., $\bm{n}_{y}\neq\emptyset$), excluding $\bm{n}$ from $\bm{z}$ will lower predictive performance.
We next examine whether $\bm{n}_{y}$ and $\bm{n}_{\bar{y}}$ are present in the representations across different CBM variants and datasets. For this purpose, we define the following task–concept configurations: (i) MPI3D[29], where $\bm{y}$ is the object shape, $\bm{n}_{y}$ the horizontal axis, $\bm{n}_{\bar{y}}$ the vertical axis, and $\bm{c}$ the remaining generative factors; (ii) Shapes3D[30], where $\bm{y}$ is the shape, $\bm{n}_{y}$ the floor color and wall color, $\bm{n}_{\bar{y}}$ the orientation, and $\bm{c}$ the remaining factors; (iii) CIFAR-10[31], where $\bm{y}$ is the standard classification task, $\bm{c}$ consists of 64 of the 143 attributes extracted by Oikarinen et al. [32] using GPT-3 [33], and $\bm{n}_{y}$ the remaining attributes, since all nuisances are correlated with $\bm{y}$; and (iv) CUB[34], where $\bm{y}$ is the bird species, $\bm{c}$ includes concepts from twelve randomly selected attribute groups, and $\bm{n}_{y}$ the attributes from the remaining 20 groups. While MCBMs natively support multiclass concepts, the other baselines in our study are limited to binary concepts. To ensure fairness, factors with $k$ classes are therefore represented as $k$ binary concepts.
Table 3: Average value of URR for task-related nuisances $\bm{n}_{y}$.
MPI3D Shapes3D CIFAR-10 CUB
Vanilla $35.0\pm 1.9$ $45.5\pm 6.4$ $19.8\pm 0.7$ $3.8\pm 1.0$
CBM [9] $28.1\pm 0.5$ $18.1\pm 2.9$ $18.5\pm 0.7$ $3.8\pm 0.8$
CEM [19] $43.2\pm 5.2$ $15.8\pm 3.9$ $27.2\pm 0.8$ $3.9\pm 1.1$
ARCBM [18] $28.2\pm 1.7$ $18.4\pm 2.1$ $18.2\pm 0.6$ $3.9\pm 0.9$
SCBM [23] $24.3\pm 0.4$ $21.8\pm 1.4$ $18.3\pm 0.7$ $3.6\pm 0.8$
MCBM (low $\gamma$) $10.7\pm 0.1$ $2.5\pm 0.1$ $18.0\pm 0.5$ $3.4\pm 0.9$
MCBM (medium $\gamma$) $6.7\pm 0.2$ $0.2\pm 0.3$ $18.1\pm 0.5$ $2.8\pm 0.8$
MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $\bm{17.6\pm 0.5}$ $\bm{2.4\pm 1.0}$
Task-related information leakage
To measure the presence of a nuisance factor $n_{j}\in\bm{n}_{y}$ in $\bm{z}$, we estimate $I(N_{j};Z\mid C)$—the information about $n_{j}$ contained in $\bm{z}$ beyond what is explained by $\bm{c}$. Since this quantity is intractable, we approximate it as $\hat{I}(N_{j};Z|C)=\hat{H}(N_{j}|C)-\hat{H}(N_{j}|C,Z)$, where $\hat{H}(N_{j}\mid C)=-\sum_{k=1}^{N}\log h_{\psi}^{c}(\bm{c}^{(k)})$ and $\hat{H}(N_{j}\mid C,Z)=-\sum_{k=1}^{N}\log h_{\psi}^{cz}(\bm{c}^{(k)},\bm{z}^{(k)})$. Here, $h_{\psi}^{c}$ and $h_{\psi}^{cz}$ are MLP classifiers trained to predict $n_{j}$ from $\bm{c}$ and $(\bm{c},\bm{z})$, respectively. In Table 3, we report the average value of $\frac{\hat{I}(N_{j};Z\mid C)}{H(N_{j})}$ across all $n_{j}\in\bm{n}_{y}$, which we call Uncertainty Reduction Ratio (URR). From these results, we conclude that: (i) CBMs tend to reduce nuisance information compared to VMs, though not consistently; (ii) CEMs generally preserve the largest amount of nuisance information, often exceeding even VMs; (iii) ARCBMs and HCBMs show no systematic advantage over CBMs in terms of nuisance removal; and (iv) MCBMs provide the strongest reduction of nuisance information, particularly as $\gamma$ increases, which enforces a stricter bottleneck.
Task-unrelated information leakage
Table 4: Average value of URR for task-unrelated nuisances.
MPI3D Shapes3D
Vanilla $11.3\pm 0.1$ $42.7\pm 9.1$
CBM $7.4\pm 1.9$ $20.6\pm 3.3$
CEM $15.5\pm 4.2$ $40.9\pm 1.8$
ARCBM $8.7\pm 1.4$ $26.6\pm 0.5$
SCBM $7.0\pm 0.3$ $21.7\pm 1.7$
MCBM (l $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
MCBM (m $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
MCBM (h $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
Neural networks typically discard input information, as their layers are non-invertible [35, 36]. Moreover, since $\bm{n}_{\bar{y}}$ is irrelevant for predicting $\bm{y}$, there is no incentive to retain it in the representation $\bm{z}$. One might therefore expect $\bm{z}$ to be free of $\bm{n}_{\bar{y}}$. However, prior work has shown that neural representations often preserve information not directly related to the task [24, 37]. To examine this, we analyze whether $\bm{n}_{\bar{y}}$ is present in $\bm{z}$ for MPI3D and Shapes3D—the only settings where $\bm{n}_{\bar{y}}$ is non-empty. Table 4 reports URR values for the nuisance variables in $\bm{n}_{\bar{y}}$ in. We find that: (i) as in Table 3, CEMs retain more nuisance information than even VMs; and (ii) MCBMs are the only models that consistently eliminate nuisances across all values of $\gamma$, as expected: with no incentive to preserve $\bm{n}_{\bar{y}}$ and an explicit penalty for doing so, such information is naturally discarded.
Are concepts removed to a greater extent in MCBMs?
Table 5: Average concepts accuracy
CIFAR-10 CUB
CBM $84.8\pm 0.2$ $96.3\pm 0.1$
CEM $84.8\pm 0.2$ $96.3\pm 0.1$
ARCBM $84.3\pm 0.2$ $96.2\pm 0.1$
SCBM $84.3\pm 0.2$ $\bm{96.5\pm 0.1}$
MCBM (l $\gamma$) $\bm{84.9\pm 0.1}$ $96.3\pm 0.2$
MCBM (m $\gamma$) $\bm{84.9\pm 0.2}$ $96.1\pm 0.1$
MCBM (h $\gamma$) $84.8\pm 0.1$ $95.8\pm 0.3$
One might worry that removing nuisance information could also inadvertently eliminate information about the concepts. To test this, we report concept prediction accuracy for CIFAR-10 and CUB in Table 5 (note that all models reach 100% accuracy on MPI3D and Shapes3D). We can observe that (i) no model consistently outperforms the others—some preserve more concept information in CIFAR-10, while others perform slightly better in CUB; and (ii) increasing $\gamma$ in MCBMs gradually reduces concept accuracy, as stronger regularization may suppress features correlated with the concepts. These results indicate that MCBMs effectively remove nuisance information while largely preserving concept-relevant content.
How does this affect task performance?
Table 6: Task accuracy
MPI3D Shapes3D CIFAR-10 CUB
Vanilla $\pm 0.0 100.0±0.010$$\pm 0.0$ $2.1\pm 0.6 77$$7.4\pm 0.3$
CBM $99.9\pm 0.0$ $100.0\pm 0.0$ $3.8\pm 0.1 77$$7.6\pm 0.2$
CEM $\pm 0.0 100.0±0.010$$\pm 0.0$ $3.1\pm 0.3 77$$7.5\pm 0.3$
ARCBM $24.2\pm 0.6$ $29.7\pm 0.4$ $68.9\pm 0.3$ $75.5\pm 0.8$
SCBM $24.2\pm 0.5$ $29.7\pm 0.3$ $62.3\pm 0.0$ $73.5\pm 0.4$
MCBM (l $\gamma$) $92.7\pm 1.1$ $100.0\pm 0.0$ $72.4\pm 0.5$ $78.3\pm 0.6$
MCBM (m $\gamma$) $46.0\pm 0.6$ $32.2\pm 1.5$ $0.8\pm 0.2 77$$7.4\pm 0.4$
MCBM (h $\gamma$) $24.9\pm 0.3$ $30.1\pm 0.2$ $70.5\pm 0.6$ $73.5\pm 1.5$
As previously discussed, when $\bm{n}_{y}$ is non-empty, restricting $\bm{z}$ to contain only information about $\bm{c}$ should reduce performance on $\bm{y}$: if $\bm{c}$ alone is insufficient to solve $\bm{y}$, then $\bm{z}$ cannot achieve perfect accuracy. In Table 6, we observe the following: (i) CBMs and CEMs reach task accuracy comparable to (or even higher than) VMs, indicating that they also rely on $\bm{n}_{y}$ to predict $\bm{y}$; (ii) ARCBMs and SCBMs achieve lower task accuracy, as they impose a bottleneck after the representations (see Appendix D); and (iii) MCBMs show decreasing task accuracy as $\gamma$ increases, reflecting the stricter bottleneck applied to the representations. Importantly, this reduction in task accuracy should not be viewed negatively: since all models achieve similar concept accuracy (see Table 5), it indicates that predictions rely less on nuisance information.
4.2 Do MCBMs Yield More Interpretable Representations?
Figure 4: CKA (left) and Disentanglement (right).
One of the key properties often attributed to CBMs is that their internal representations align with human-interpretable concepts. While this generally holds, we show that MCBMs yield representations that are even more interpretable. To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) [38, 39, 40], which measures the similarity between learned representations and concept labels (encoded as one-hot vectors); and (ii) Disentanglement[41], which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$. These metrics capture interpretability from complementary perspectives: (i) alignment indicates whether the information is organized in a concept-aware fashion; and (ii) disentanglement reflects the extent to which individual concepts are independently encoded in the representations. As shown in Figure 4: (i) CBMs, CEMs, ARCBMs, and HCBMs achieve stronger alignment with concepts than Vanilla Models, yet they do not consistently improve disentanglement; and (ii) MCBMs consistently improve both alignment and disentanglement, particularly as $\gamma$ increases, reflecting the removal of additional nuisances. These results suggest that explicitly eliminating nuisances leads to more interpretable representations.
4.3 Do MCBMs Enable More Reliable Interventions?
(a) CIFAR-10
(b) CUB
Figure 5: Error (y-axis) versus percentage of concepts intervened (x-axis) across different models.
Another key property often attributed to CBMs is their capacity to support interventions. Although standard CBM intervention methods suffer from theoretical limitations (see Section 5), they can still be applied in practice. Accordingly, for CBMs we follow the procedure in Section 2.3, for ARCBMs and SCBMs we use the procedure described in Appendix D, and for MCBMs we apply the method in Section 2.4. To evaluate intervention reliability across models, we adopt the standardized protocol of Koh et al. [9], which tracks how prediction error changes as the number of intervened concepts increases. Following Shin et al. [42], we intervene on the concepts predicted with the lowest confidence. Results for CIFAR-10 and CUB are shown in Figure 5 (note that interventions have minimal effect on MPI3D and Shapes3D, since—as shown in Table 6—the concepts are not strong predictors of the task). From these results, four main observations emerge: (i) CBMs may even increase error when multiple concepts are intervened—an effect of nuisance information leaking into the representation due to the absence of a proper bottleneck; (ii) ARCBMs and SCBMs mitigate this by applying an Information Bottleneck after the representations, which improves intervention effectiveness while leaving interpretability unchanged (Figure 4); (iii) in MCBMs, intervention gains remain largely invariant to $\gamma$, as indicated by the nearly parallel trends across its different values; and (iv) at low and medium $\gamma$, MCBMs deliver the strongest intervention performance, especially when fewer than 100% of the concepts are intervened, clearly outperforming ARCBMs and SCBMs.
5 Other Fundamental Theoretical Flaws of CBMs
As briefly discussed in Section 2.3, CBMs—unlike MCBMs—do not provide a principled mechanism to estimate $p(z_{j}\mid c_{j})$. To enable interventions, two assumptions are typically introduced: (i) multiclass concepts are handled using a One-vs-Rest scheme, and (ii) interventions are implemented via the sigmoid inverse function, i.e., $p(z_{j}\mid c_{j})\approx\delta\left(z_{j}-\sigma^{-1}(c_{j})\right)$. These assumptions are not only ad hoc but also theoretically incorrect, as we explain and illustrate with toy experiments below.
One-vs-Rest Limitations
Figure 6: CBM (top) versus MCBM (bottom): class boundaries (left), highest predicted likelihood (middle), and second-highest predicted likelihood (right).
One-vs-Rest strategies exhibit several limitations: (i) individual binary classifiers tend to be biased toward the negative class, and (ii) the predicted probabilities across classifiers are typically uncalibrated [43]. To illustrate these issues, we design an experiment where the concepts are defined based on a four-class spiral dataset with imbalanced class distributions. We train: (i) a CBM with One-vs-Rest binarization, and (ii) an MCBM modeling concepts directly as multiclass variables.
The results reveal three major shortcomings of CBMs trained with the One-vs-Rest strategy. First, Figure 6 (left) shows that they overpredict the most frequent class, especially when the true class is rare and spatially close to the dominant. Second, Figure 6 (middle) shows that CBMs produce poorly calibrated predictions, lacking the smooth likelihood transitions observed in MCBMs. Finally, Figure 6 (right) shows that CBMs often assign near-one likelihoods to multiple classes simultaneously, whereas MCBMs confine high secondary likelihoods to regions of overlap, producing more reliable uncertainty estimates. Importantly, these effects may not be captured by standard metrics like accuracy, yet they represent fundamental flaws from a probabilistic perspective.
Sigmoid Inverse Function
Figure 7: Density (y-axis) vs. $z$ (x-axis)
[…]
6 Conclusions
In this paper, we argue that—contrary to common belief—Concept Bottleneck Models (CBMs) do not enforce a true bottleneck: although representations are encouraged to retain concept-related information, they are not constrained to discard nuisance information. This limitation undermines interpretability and provides no theoretical guarantees for intervention procedures. To address this, we propose Minimal Concept Bottleneck Models (MCBMs), which introduce an Information Bottleneck in the representation space via an additional loss term derived through variational approximations. Beyond enforcing a proper bottleneck, our formulation ensures that interventions remain consistent with Bayesian principles. Empirically, we show that CBMs and their variants fail to remove non-concept information from the representation, even when irrelevant to the target task. In contrast, MCBMs effectively eliminate such information while preserving concept-relevant content, yielding more interpretable representations and principled interventions. Finally, we highlight fundamental flaws in the intervention process of CBMs, which MCBMs overcome due to their principled design.
Acknowledgments and Disclosure of Funding
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007666, MCIN/AEI/10.13039/501100011033 under Grant PID2021-126061OB-C44, and the Government of Aragón (Grant Group T36 23R).
Appendix A Complete Graphical Models
(a) Data Generative Process
(b) Vanilla Model
(c) Concept Bottleneck Models
(d) Minimal Concept Bottleneck Models (ours)
Figure 8: Graphical models of the different systems described. We consider $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $\color[rgb]{0,1,1}p_{\theta}(\bm{z}|\bm{x})$ (cyan arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}|\bm{z})$ (green arrows). Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q(z_{j}|c_{j})$ (red arrows).
Appendix B Details of Derivations
B.1 Proof of Equation 1
$I(Z;Y)$ $=\iint p(z,y)\log\frac{p(y|z)}{p(y)}\,dy\,dz$ (7)
$=\iiint p(x,y)p_{\theta}(z|x)\log\frac{p(y|z)}{p(y)}\,dx\,dy\,dz$ (8)
$=\iiint p(x,y)p_{\theta}(z|x)\log\frac{p(y|z)}{p(y)}\frac{q_{\phi}(\hat{y}|z)}{q_{\phi}(\hat{y}|z)}\,dx\,dy\,dz$ (9)
$=\mathbb{E}_{p(x,y)}\left[\mathbb{E}_{p_{\theta}(z|x)}\left[\log q_{\phi}(\hat{y}|z)\right]\right]+E_{p_{\theta}(z)}\left[D_{KL}\left(p(y|z)||q_{\phi}(\hat{y}|z)\right)\right]+H(Y)$ (10)
$\geq\mathbb{E}_{p(x,y)}\left[\mathbb{E}_{p_{\theta}(z|x)}\left[\log q_{\phi}(\hat{y}|z)\right]\right]+H(Y)$ (11)
Thus, since $H(Y)$ is independent of $\theta$ and $\phi$, we have that:
$\max_{Z}I(Z;Y)=\max_{\theta,\phi}\mathbb{E}_{p(x,y)}\left[\mathbb{E}_{p_{\theta}(z|x)}\left[\log q_{\phi}(\hat{y}|z)\right]\right]$ (12)
B.2 Proof of Equation 2
$I(Z_{j};C_{j})$ $=\iint p(z_{j},c_{j})\log\frac{p(c_{j}|z_{j})}{p(c_{j})}\,dc_{j}\,dz_{j}$ (13)
$=\iiint p(x,c_{j})p_{\theta}(z_{j}|x)\log\frac{p(c_{j}|z_{j})}{p(c_{j})}\,dx\,dc_{j}\,dz_{j}$ (14)
$=\iiint p(x,c_{j})p_{\theta}(z_{j}|x)\log\frac{p(c_{j}|z_{j})}{p(c_{j})}\frac{q(\hat{c}_{j}|z_{j})}{q(\hat{c}_{j}|z_{j})}\,dx\,dc_{j}\,dz_{j}$ (15)
$\geq E_{p(x,c_{j})}\left[E_{p_{\theta}(z_{j}|x)}\left[\log q(\hat{c}_{j}|z_{j})\right]\right]+E_{p_{\theta}(z_{j})}\left[D_{KL}\left(p(c_{j}|z_{j})||q(\hat{c}_{j}|z_{j})\right)\right]+H(C_{j})$ (16)
$\geq E_{p(x,c_{j})}\left[E_{p_{\theta}(z_{j}|x)}\left[\log q(\hat{c}_{j}|z_{j})\right]\right]+H(C_{j})$ (17)
Given the fact that $H(C_{j})$ is constant, we have that:
$\max_{Z_{j}}I(Z_{j};C_{j})=\max_{\theta}E_{p(x,c_{j})}\left[E_{p_{\theta}(z_{j}|x)}\left[\log q(\hat{c}_{j}|z_{j})\right]\right]$ (18)
B.3 Proof of Equation 4
$I(Z_{j};X|C_{j})$ $=\iiint p(x,c_{j})p_{\theta}(z_{j}|x)\log\frac{p(z_{j}|x)}{p(z_{j}|c_{j})}\,dx\,dc_{j}\,dz_{j}$ (19)
$=\iiint p(x,c_{j})p_{\theta}(z_{j}|x)\log\frac{p(z_{j}|x)}{p(z_{j}|c_{j})}\frac{q(\hat{z}_{j}|c_{j})}{q(\hat{z}_{j}|c_{j})}\,dx\,dc_{j}\,dz_{j}$ (20)
$=\mathbb{E}_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta}(z_{j}|x)||q(\hat{z}_{j}|c_{j})\right)\right]-E_{p(c_{j})}\left[D_{KL}\left(p(z_{j}|c_{j})||q(\hat{z}_{j}|c_{j})\right)\right]$ (21)
$\leq\mathbb{E}_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta}(z_{j}|x)||q(\hat{z}_{j}|c_{j})\right)\right]$ (22)
Thus, we have that:
$\min_{Z_{j}}I(Z_{j};X|C_{j})=\min_{\theta}E_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta}(z_{j}|x)||q(\hat{z}_{j}|c_{j})\right)\right]$ (23)
B.4 KL Divergence between two Gaussian Distributions
We are given the conditional distributions:
$p_{\theta}(z_{j}|x)$ $=\mathcal{N}(f_{\theta}(x)_{j},\sigma_{x}^{2}I),$
$q_{\phi}(\hat{z}_{j}|c_{j})$ $=\mathcal{N}(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I),$
and aim to minimize the expected KL divergence:
$\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[D_{\text{KL}}\left(p_{\theta}(z_{j}|x)\;\|\;q_{\phi}(\hat{z}_{j}|c_{j})\right)\right].$
The KL divergence between two multivariate Gaussians with diagonal covariances is given by:
$D_{\text{KL}}(\mathcal{N}(\mu_{p},\Sigma_{p})\;\|\;\mathcal{N}(\mu_{q},\Sigma_{q}))=\frac{1}{2}\left[\operatorname{tr}(\Sigma_{q}^{-1}\Sigma_{p})+(\mu_{q}-\mu_{p})^{\top}\Sigma_{q}^{-1}(\mu_{q}-\mu_{p})-d+\log\frac{\det\Sigma_{q}}{\det\Sigma_{p}}\right].$
Applying this to our case:
  • $\mu_{p}=f_{\theta}(x)_{j}$, $\mu_{q}=g^{z}_{\phi}(c_{j})$,
  • $\Sigma_{p}=\sigma_{x}^{2}I$, $\Sigma_{q}=\sigma_{\hat{z}}^{2}I$,
  • $d$ is the dimension of $z_{j}$.
Plugging in, we obtain:
$D_{\text{KL}}=\frac{1}{2}\left[\frac{d\sigma_{x}^{2}}{\sigma_{\hat{z}}^{2}}+\frac{1}{\sigma_{\hat{z}}^{2}}\|f_{\theta}(x)_{j}-g^{z}_{\phi}(c_{j})\|^{2}-d+d\log\left(\frac{\sigma_{\hat{z}}^{2}}{\sigma_{x}^{2}}\right)\right].$
Note that all terms except the squared distance are constant with respect to $\theta$ and $\phi$. Therefore:
$\mathbb{E}_{p(x,c_{j})}\left[D_{\text{KL}}\left(p_{\theta}(z_{j}|x)\;\|\;q_{\phi}(\hat{z}_{j}|c_{j})\right)\right]=\frac{1}{2\sigma_{\hat{z}}^{2}}\mathbb{E}_{p(x,c_{j})}\left[\|f_{\theta}(x)_{j}-g^{z}_{\phi}(c_{j})\|^{2}\right]+\text{const}.$
Conclusion: Minimizing the expected KL divergence
$\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[D_{\text{KL}}\left(p_{\theta}(z_{j}|x)\;\|\;q_{\phi}(\hat{z}_{j}|c_{j})\right)\right]$
is equivalent (up to a scaling factor) to minimizing the expected mean squared error:
$\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[\|f_{\theta}(x)_{j}-g^{z}_{\phi}(c_{j})\|^{2}\right].$
Appendix C Training Algorithm of MCBMs
Algorithm 1 Training Algorithm for MCBMs
1:Dataset $\mathcal{D}=\{\bm{x}^{(k)},\bm{y}^{(k)},\bm{c}^{(k)}\}_{k=1}^{N}$, latent norm $\lambda$, learning rate $\eta$, batch size $B$
2:Parameters $\theta$ (encoder), $\phi$ (class-head, task-heads, representation-heads)
3:Initialize parameters $\theta$, $\phi$ and representations heads:
4:for all $j=1,\dots,n$ do
5:  if $c_{j}$ is binary then
6:   $g^{z}_{j}\leftarrow\lambda\text{ if }c_{j}=1\text{ else }-\lambda$
7:  else if $c_{j}$ is multiclass then
8:   $g^{z}_{j}\leftarrow\lambda\cdot\mathrm{one\_hot}\left(c_{j}\right)$
9:  else
10:   $g^{z}_{j}\leftarrow\lambda\cdot c_{j}$
11:  end if
12:end for
13:while not converged do
14:  Sample a mini-batch $\{\bm{x}^{(k)},\bm{y}^{(k)},\bm{c}^{(k)}\}_{k=1}^{B}\sim\mathcal{D}$
15:  for all $\bm{x}^{(k)},\bm{y}^{(k)},\bm{c}^{(k)}$ in batch do
16:   Encode: Compute $\mu_{\theta}^{(k)}\leftarrow f_{\theta}\left(x^{(k)}\right)$
17:   Sample noise $\epsilon\sim\mathcal{N}(0,I)$ $\triangleright$ Reparameterization trick with only one sample
18:   Reparameterize: $\bm{z}^{(k)}\leftarrow\mu_{\theta}^{(k)}+\sigma_{x}\odot\epsilon$
19:   Task prediction: $\hat{\bm{y}}^{(k)}\leftarrow g^{y}_{\phi}\left(\bm{z}^{(k)}\right)$ $\triangleright$ Similar to VMs
20:   Task loss: $\mathcal{L}_{y}^{(k)}\leftarrow\left\|\bm{y}^{(k)}-\hat{\bm{y}}^{(k)}\right\|^{2}\text{ if }\bm{y}\text{ is continuous else }\text{CE}\left(\bm{y}^{(k)},\bm{\hat{y}}^{(k)}\right)$
21:   for all $j=1,\dots,n$ do
22:     Concept $j$ prediction: $\hat{c}^{(k)}_{j}\leftarrow g^{c}_{\phi,j}\left(z^{(k)}_{j}\right) \triangleright$ Similar to CBMs
23:     Concept $j$ loss: $\mathcal{L}_{c,j}^{(k)}\leftarrow\left\|c^{(k)}_{j}-\hat{c}^{(k)}_{j}\right\|^{2}\text{ if }c_{j}\text{ is continuous else }\text{CE}\left(c^{(k)}_{j},\hat{c}^{(k)}_{j}\right)$
24:     Representation $j$ prediction: $\hat{z}^{(k)}_{j}\leftarrow g^{z}_{j}\left(c^{(k)}_{j}\right) \triangleright$ Novelty in MCBMs
25:     Representation $j$ loss: $\mathcal{L}_{z,j}^{(k)}\leftarrow\left\|z^{(k)}_{j}-\hat{z}^{(k)}_{j}\right\|^{2}$
26:   end for
27:   Total loss: $\mathcal{L}^{(k)}\leftarrow\mathcal{L}_{y}^{(k)}+\beta\sum_{j=1}^{n}\mathcal{L}_{c,j}^{(k)}+\gamma\sum_{j=1}^{n}\mathcal{L}_{z,j}^{(k)}$
28:  end for
29:  Update $\theta,\phi$ using gradient descent:
$\theta\leftarrow\theta-\eta\nabla_{\theta}\left(\frac{1}{B}\sum_{k=1}^{B}\mathcal{L}^{(k)}\right),\quad\phi\leftarrow\phi-\eta\nabla_{\phi}\left(\frac{1}{B}\sum_{k=1}^{B}\mathcal{L}^{(k)}\right)$
30:end while
Appendix D Hard Concept Bottleneck Models
As discussed in Section 3, one of the most successful approaches to mitigating information leakage is the family of Hard Concept Bottleneck Models (HCBMs) [18]. Unlike MCBMs, which enforce an Information Bottleneck (IB) directly at the representation level $\bm{z}$, HCBMs operate by pruning the information contained in $\bm{z}$ before producing the task prediction $\hat{\bm{y}}$. In doing so, they encourage $\hat{\bm{y}}$ to depend solely on $\bm{c}$, thereby yielding interventions that are more reliable in practice. This approach differs fundamentally from MCBMs: whereas Equation 4 in MCBMs minimizes over $Z_{j}$, thereby removing nuisances directly from the representation $z_{j}$, HCBMs optimize at the prediction level, focusing on $\hat{Y}$. Formally, they are trained to minimize:
$\vskip 2.0pt\min_{\hat{Y}}I(\hat{Y};X|C)=\min_{\theta,\phi}\mathbb{E}_{p(x,c_{j})}\left[D_{KL}\left(p_{\theta,\phi}(\hat{\bm{y}}|x)||p(\hat{\bm{y}}|c)\right)\right]$ (24)
That is, the optimization is performed directly over the predictions $\hat{\bm{y}}$. To implement this, it defines:
$p_{\theta,\phi}(\hat{\bm{y}}|x)=\iiint q_{\phi}\left(\hat{\bm{y}}|\hat{\bm{c}}^{b}\right){\color[rgb]{1,0,1}p\left(\hat{\bm{c}}^{b}|\hat{\bm{c}}\right)}q\left(\hat{\bm{c}}|\bm{z}\right)p_{\theta}\left(\bm{z}|\bm{x}\right)\,d\hat{\bm{c}}^{b}\,d\hat{\bm{c}}\,d\bm{z}$ (25)
where $q_{\phi}\left(\hat{\bm{y}}\mid\hat{\bm{c}}^{b}\right)$ denotes the new task head, $q\left(\hat{\bm{c}}\mid\bm{z}\right)$ the concept head, and $p_{\theta}\left(\bm{z}\mid\bm{x}\right)$ the encoder, as defined in Section 2. The distribution
$p\left(\hat{c}_{j}^{b}|\hat{c}_{j}\right)=\delta\left(\hat{c}_{j}^{b}-\Theta\left(\hat{c}_{j}-0.5\right)\right)$ (26)
referred to as the binarizing head, applies the Heaviside step function $\Theta$ to produce a binary version $\hat{c}^{b}_{j}\in\{0,1\}$ of $\hat{c}_{j}$. In this way, HCBMs enforce an ad-hoc Information Bottleneck by predicting $\hat{\bm{y}}$ from binarized concept representations. This process is schematized in Figure 9.
How are Interventions Performed in HCBMs?
Because HCBMs introduce a bottleneck immediately before predicting $\hat{\bm{y}}$, the intervention process is more straightforward than in standard CBMs. More specifically, interventions are carried out according to:
$p(\hat{\bm{y}}|c_{j}=\alpha,\bm{x})=\iint q_{\phi}(\hat{\bm{y}}|\hat{c}^{b}_{j},\bm{\hat{c}}^{b}_{\backslash j})p(\hat{c}^{b}_{j}|c_{j}=\alpha)p_{\theta}(\bm{\hat{c}}^{b}_{\backslash j}|\bm{x})\,d\hat{c}^{b}_{j}\,d\bm{\hat{c}}^{b}_{\backslash j}$ (27)
Here, $p(\hat{c}^{b}_{j}|c_{j}=\alpha)$ is approximated as $\delta(\hat{c}^{b}_{j}-\alpha)$, while the other distributions are available in closed form. Although this procedure provides stronger guarantees than the intervention mechanisms in standard CBMs, two main issues remain:
  1. (i)
    The optimization is performed over the predictions $\hat{\bm{y}}$ instead of the representation $\bm{z}$. As a result, the representations themselves are not necessarily interpretable, as evidenced in Figure 4, which limits one of the core motivations for adopting CBMs in the first place.
  2. (ii)
    HCBMs still require binarizing multiclass concepts, which introduces theoretical limitations and practical drawbacks, as further discussed in Section 5.
Figure 9: Graphical models of HCBMs with two concepts and two-dimensional representations. Hard Concept Bottleneck Models obtain a binarized version $\hat{c}^{b}_{j}$ of each predicted concept $\hat{c}_{j}$ through the binarizing head $\color[rgb]{1,0,1}p(\hat{c}^{b}_{j}\mid\hat{c}j)$ (fuchsia arrows). Unlike the models in Figure 2, HCBMs predict the task output $\hat{\bm{y}}$ from the binarized concepts $\hat{c}^{b}_{j}$ using the new task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ (green arrows).
Appendix E Experiments Details
E.1 Hyperparameters for Section 4
Table 7: Hyperparameters for Section 4. For all datasets, the concept head $g^{c}_{\phi}$ is implemented as the identity function in CBMs, and as a multilayer perceptron (MLP) with three hidden layers in MCBMs. We emphasize that CBMs are, by design, restricted to use invertible $g^{c}_{\phi}$ to enable intervention procedures.
MPI3D Shapes3D CIFAR-10 CUB
$f_{\theta}$ architecture ResNet20 ResNet20 2 conv. layers InceptionV3
$f_{\theta}$ pretraining None None None ImageNet
$g^{y}_{\phi}$ hidden layers 64 64 64 256
low $\gamma$ 1 1 0.1 0.05
medium $\gamma$ 3 3 0.3 0.1
high $\gamma$ 5 5 0.5 0.3
number of epochs 50 50 200 250
batch size 128 128 128 128
optimizer SGD SGD Adam SGD
learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $1\times 10^{-4}$ $2\times 10^{-2}$
momentum 0.9 0.9 0. 0.9
weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
scheduler Step Step Step Step
step size (epochs) 20 20 80 100
scheduler $\gamma$ 0.1 0.1 0.1 0.1
E.2 Datasets
MPI3D
This is a synthetic dataset with controlled variation across seven generative factors: object shape, object color, object size, camera height, background color, horizontal axis, and vertical axis. In our setup, $\bm{y}$ corresponds to the object shape, $\bm{n}_{y}$ to the horizontal axis, $\bm{n}_{\bar{y}}$ to the vertical axis, and $\bm{c}$ to the remaining generative factors. To ensure consistency in the mapping between concepts and task nuisances and the target, we filter the dataset such that any combination of elements in $\left\{\bm{c},\bm{n}_{y}\right\}$ corresponds to a unique value of $\bm{y}$. All invalid combinations are removed accordingly.
Shapes3D
This synthetic dataset consists of 3D-rendered objects placed in a room, with variation across six known generative factors: floor color, wall color, object color, scale, shape, and orientation. In our setup, $\bm{y}$ corresponds to the shape, $\bm{n}y$ includes floor color and wall color, $\bm{n}{\bar{y}}$ corresponds to orientation, and $\bm{c}$ comprises the remaining factors. We follow the same filtering strategy as in MPI3D to construct this configuration: we retain only those samples for which each combination of $\left\{\bm{c},\bm{n}_{y}\right\}$ uniquely determines $\bm{y}$, removing all invalid configurations.
CIFAR-10
CIFAR-10 is a widely used image classification benchmark consisting of 60,000 natural images of size 32 $\times$ 32, divided into 10 classes (e.g., airplanes, automobiles, birds, cats, etc.). The dataset is split into 50,000 training and 10,000 test images, with balanced class distributions. To reduce the need for manual concept annotations, the concepts are synthetically derived following the methodology of [23]. A total of 143 attributes are extracted using GPT-3 [33]; 64 form the concept set $\bm{c}$, while the rest define the nuisance set $\bm{n}_{y}$. Binary values are obtained with the CLIP model [44] by comparing the similarity of each image to the embedding of an attribute and to its negative counterpart.
CUB
The Caltech-UCSD Birds (CUB) dataset contains 11,788 images of 200 bird species, annotated with part locations, bounding boxes, and 312 binary attributes. Following the approach of Koh et al. [9], we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes. These attributes are grouped into 27 semantic clusters, where each group is defined by a common prefix in the attribute names. In our setup, the task variable $\bm{y}$ is to classify the bird species. The concept set $\bm{c}$ consists of the attributes belonging to 12 randomly selected groups (per run), while the nuisance set $\bm{n}y$ includes the attributes from the remaining 20 groups. Since most attributes exhibit some correlation with the classification task, we set $\bm{n}_{\bar{y}}$ to the empty set.